OpenStudy (anonymous):

find t12 for a geometric sequence where t1=2+2i and r=3

OpenStudy (jdoe0001):

did you say \(\Large t_1 = 2+2i?\)

OpenStudy (jdoe0001):

\(\large { a_{\color{brown}{ n}}=a_1\cdot r^{{\color{brown}{ n}}-1}\qquad \qquad \begin{cases} a_1\to first\ term\\ r=common\ ratio \end{cases}}\)

OpenStudy (jdoe0001):

thus \(\large { a_{\color{brown}{ 12}}=a_1\cdot r^{{\color{brown}{ 12}}-1}\qquad \qquad \begin{cases} a_1\to &first\ term\\ r\to &common\ ratio \end{cases} }\)

OpenStudy (mathmale):

Unfortunately, your presentation of this problem is hard to follow. "find t12 for a geometric sequence where t1=2+2i and r=3" What does the letter 't' represent here? t1=2+2i does not register. Where did that come from? Note that there is no addition or subtraction in a geometric sequence. So t1=2+2i could not be right. Type in your question again. This time, please use real subscripts for subscripts: